# Physics beyond the Standard Model with the DSA-2000

### v1

@@ 4 Pulsar timing arrays
Millisecond pulsars (MSPs) have long been used as detectors of nanoHertz gravitational waves (GWs) [168]. By measuring the timing of radio pulses from an array of MSPs accurately, one can measure ripples of spacetime manifesting as deviations in pulsar timings. In June 2023, The North American Nanohertz Observatory for Gravitational Waves (NANOGrav) announced positive evidence for a $\sim 3\sigma$ evidence of the existence of a stochastic gravitational wave background, possibly due to a population of merging supermassive BHs [169]. The DSA-2000 is projected to detect approximately 3,000 millisecond pulsars [170], and to increase the total number of high-quality pulsars being timed to 200, with a monthly cadence [1], significantly improving NANOGrav’s sensitivity to gravitational waves.
In addition to gravitational waves, PTAs have also been proposed as probes of DM substructure and primordial black holes (PBHs) [171, 172, 173, 174, 175, 176, 177, 178, 179, 180, 30]. DM substructure are self-gravitating DM objects formed on small ($\lesssim$ kpc) scales. While the large-scale matter power spectrum is accurately measured from the cosmic microwave background (CMB), DM substructure on small scales is much more difficult to detect. More importantly, different DM theories can predict different small-scale structures [181, 182, 183, 184, 185], making the detection of these objects an important step in understanding the nature of DM. In this section, we review the idea of detecting DM substructure using PTAs, and explain how to calculate the projected sensitivity for a future PTA survey, including timing data acquired by DSA-2000. We find that the detection of DM substructure may be possible with sufficient observation time.
@@ 4.1 Gravitational effects
[…]
where the timing residuals, $h(t)$, are defined to be the measured pulsar time shifts subtracting the best-fitted pulsar timing evolution consistent with the timing model of a naturally evolving pulsar [168],
@@
where $\mathbf{d}$ is the vector along the line-of-sight between the pulsar and Earth. Combining Eqs. (4.1)-(4.4), one obtains the timing residual due to the Doppler effect from a transiting DM substructure [177, 186],
$h(t)=\frac{GM}{v^{2}}\mathbf{\hat{d}}\cdot\left\{\mathbf{\hat{b}}\sqrt{1+\left[\frac{v(t-t_{0})}{b}\right]^{2}}-\mathbf{\hat{v}}\sinh^{-1}\left[\frac{v(t-t_{0})}{b}\right]\right\}-\delta t_{\mathrm{fitted}}(t)\,.$ (4.5)
Here we have absorbed all terms up to $\mathcal{O}(t^{2})$ into $\delta t_{\mathrm{fitted}}(t)$, since these terms are typically included in a fit within the timing model. Practically, when searching for physics signals within the PTA dataset, the fitting part is handled by software packages such as ENTERPRISE [187], ENTERPRISE_EXTENSIONS [188], and a recently developed wrapper, PTArcade [189]. See Ref. [190] for the Bayesian inference framework used to search for DM substructure with PTA data. Hence, the $\delta t_{\mathrm{fitted}}(t)$ term in Eq. (4.5) can generally be neglected when the signal shape is interfaced into these programs for searching the actual dataset. However, for the purpose of making sensitivity projections for future experiments, the $\delta t_{\mathrm{fitted}}(t)$ term can lead to a reduction in sensitivity if a portion of the signal is degenerate with the timing model and gets absorbed into the fit [179].
Given a total observation time of $T$ on a particular pulsar, in the limit where $b\gg vT$ (which is applicable for DM masses sufficiently large), Eq. (4.5) can be expanded in a power series in $t$ [^11] One can also consider signals in the opposite limit (i.e., the dynamic limit) and effects such as the Shapiro delay. However, their upper limits are generally weaker than that of the Doppler dynamic effect when analyzing real data [30], so we do not include them in this work. The terms up to quadratic order in $t$ are degenerate with the intrinsic pulsar timing model and will be absorbed into the fit. Therefore, we parameterize the timing residual starting from the cubic order in $t$ [186, 190],
$h(t)=\frac{A}{\mathrm{yr}^{2}}t^{3}\,,$ (4.6)
@@
[…]
@@
$\mathrm{SNR}=\frac{GMT^{7/2}v|\mathbf{\hat{v}}\cdot\mathbf{\hat{d}}|}{48\sqrt{7}b^{3}\Sigma}\,.$ (4.8)
@@
The above analysis considers the SNR from timing measurements of a single pulsar. We now generalize this to the case of PTAs, where timing measurements from multiple pulsars are available. Since the DM phase space is a statistical distribution near each pulsar, the expected SNR from each pulsar also follows a statistical distribution. The variance in DM velocity can generally be neglected in an order-of-magnitude estimate due to its linear scaling in the SNR. Statistically, the signal is dominated by the DM with the smallest impact parameter. For a given pulsar, this can be derived by considering putting random points in a spherical volume, in which the cumulative distribution function (CDF) of the impact parameter of each DM is given by $F_{B}(b)=(b/R)^{3}$, where $R$ is the radius of a large (fictitious) spherical volume, and thus the CDF of the minimal impact parameter is [177]
@@
$=\exp\left(-\frac{4\pi}{3}\frac{G\rho_{\mathrm{DM}}f_{\mathrm{DM}}}{48\sqrt{7}}\frac{1}{\mathrm{SNR}}\frac{T^{7/2}v|\mathbf{\hat{v}}\cdot\mathbf{\hat{d}}|}{\Sigma}\right)\,.$ (4.10)
@@
to be the fraction of dark matter in the form of compact objects, which are the components that can contribute to PTA signals. The theoretically compelling parameter space is $f_{\mathrm{DM}}\leq 1$, although this assumes the robustness of the canonical local dark matter density value, $\rho_{\mathrm{DM}}=0.4~{}\mathrm{GeV/cm}^{3}$. Current experimental constraints, however, allow values as large as $f_{\mathrm{DM}}\lesssim 2\times 10^{3}$, based on limits from solar system ephemerides [191], since there is currently no direct local observational measurement of $\rho_{\mathrm{DM}}$.
If we are now given timing data from $N_{P}$ pulsars, each labeled with subscript $i$, the optimal sensitivity to compact DM is achieved by considering the statistics of the pulsar with the highest SNR [190], denoted as $\mathrm{SNR}_{\max}$, whose CDF can be readily derived from Eq. (4.1), and the fact that the SNRi from each pulsar is independent:
@@
$=\exp\left(-\frac{4\pi}{3}\frac{G\rho_{\mathrm{DM}}f_{\mathrm{DM}}}{48\sqrt{7}}\frac{1}{\mathrm{SNR}_{\max}}\sum_{i=1}^{N_{P}}\frac{T_{i}^{7/2}v_{i}|\mathbf{\hat{v}}_{i}\cdot\mathbf{\hat{d}}_{i}|}{\Sigma_{i}}\right)\,.$ (4.12)
@@
Using Eq. (4.1), for any $0\leq p\leq 1$, we can derive the $100p^{\mathrm{th}}$-percentile $\mathrm{SNR}_{\max}$, denoted as $\mathrm{SNR}_{\max,p}$, by setting $F_{\mathrm{SNR}_{\max}}(\mathrm{SNR}_{\max,p})=p$,
$\mathrm{SNR}_{\max,p}$ $=\frac{4\pi}{3}\frac{G\rho_{\mathrm{DM}}f_{\mathrm{DM}}}{48\sqrt{7}}\frac{1}{\log p^{-1}}\sum_{i=1}^{N_{P}}\frac{T_{i}^{7/2}v_{i}|\mathbf{\hat{v}}_{i}\cdot\mathbf{\hat{d}}_{i}|}{\sigma_{i}}$
$\propto f_{\mathrm{DM}}\sum_{i=1}^{N_{P}}T_{i}^{7/2}\,,$ (4.13)
@@
where $N_{P}$ is the number of pulsars used in the analysis. We have assumed that the noise level for timing data from each pulsar is identical, and that the noise level for new pulsars to be discovered and timed by DSA-2000 is the same as that of existing pulsars timed by NANOGrav. Additionally, in the second line, we have neglected the variance in DM velocity and $\mathcal{O}(1)$ angular factors. Note in particular that the pulsar observing time $T_{i}$ has a particularly strong impact on the SNR, since a longer observation time leads to a smaller expectation value of $b_{\min}$, and thus a stronger signal.
@@
Figure 10: Upper limits at the 95% credible level on the local PBH abundance (left panel) and fifth force coupling strength (right panel). The blue line is the existing constraint published by the NANOGrav collaboration [30], while the orange and green lines are the projected sensitivity by 2033 and 2053, corresponding to 5 and 25 years after DSA-2000 begins taking timing data, obtained by rescaling the blue line using the projected parameters for a combined NANOGrav 25-yr / DSA 5-yr and NANOGrav 45-yr / DSA 25-yr observation dataset, and Eq. (4.14) and Eq. (4.17) (see main text). Following Ref. [30], existing constraints on $\tilde{\alpha}$ are also shown in the right panel, including bounds from the weak equivalence principle (WEP) [192, 193, 194], NS kinetic heating (assuming an additional short-range DM–baryon interaction) due to DM capture [180], and bounds derived from a combination of fifth-force constraints on baryon–baryon interactions [195, 196] and Bullet Cluster (BC) limits on DM self-interactions [197, 198] (see also Refs. [199, 180]). We note that while the BC bound is very competitive, it is an indirect constraint on $\tilde{\alpha}$ and does not apply if only a subcomponent of DM interacts through the fifth force, where, in contrast, the other constraints degrade only linearly with the DM subcomponent fraction.
@@
Figure 11: Observation times for each pulsar considered in the analysis are shown for three benchmark years: 2023 (blue), 2033 (orange), and 2053 (green). The 68 existing pulsars being timed by NANOGrav in their 15-year data release (2023) are shown at the top [200], although J0614-3329 (marked by an asterisk) is excluded from their DM substructure analysis since its observation time has been less than three years [30]. The new pulsars that DSA-2000 is expected to discover are shown at the bottom, labeled with placeholder pulsar names starting with “New Pulsar”.
@@
$\frac{f_{\mathrm{DM}}}{f_{\mathrm{DM}}^{(\mathrm{NG-15})}}=\frac{\sum_{j}^{N_{P}^{\mathrm{NG-15}}}(T^{(\mathrm{NG-15})}_{j})^{7/2}}{\sum_{i}^{N_{P}}T_{i}^{7/2}}\,,$ (4.14)
where $N_{P}^{(\mathrm{NG-15})}=67$ is the number of pulsars used in the NANOGrav 15-yr new physics analysis (one pulsar, J0614-3329, was excluded from the analysis since it was observed for less than three years) [30], released in 2023.
Note that the derivation of Eq. (4.1) assumes that only uncorrelated white noise is present in the timing data. In the presence of colored and correlated noise, the SNR from potential DM sources will be diminished, as shown in Ref. [190]. However, we expect this drop in sensitivity will be common in both the NANOGrav 15-yr dataset analysis, and any future analyses combining the NANOGrav and DSA-2000 datasets. Hence, its effect should be minimal in the projection obtained by a rescaling of the upper limit.
The above analysis shows that the sensitivity to DM substructure can be improved by (i) obtaining timing data from more pulsars, and (ii) a longer cumulative observation time for each pulsar. Both of these conditions can be achieved by utilizing DSA-2000’s power in discovering new pulsars. DSA-2000 is expected to bring the total number of pulsars being timed within the NANOGrav Collaboration to 200. In the left panel of Fig. 10, we show the projected sensitivity from the outlined procedure, assuming that the new pulsars start being timed by the collaboration in 2028. We show the projected sensitivity by 2033 and 2053, corresponding to 5 and 25 years after DSA-2000 begins taking timing data. We find that with the combined data from the NANOGrav 25-yr and DSA-2000 5-year datasets, which is the currently planned timeframe for DSA-2000, the upper limit that one can place on the local DM density in the form of compact objects can improve by roughly an order of magnitude, to $f_{\mathrm{DM}}\lesssim 40$ for $M\gtrsim 10^{-5}$ $M_{\odot}$. If we consider the combined data from the NANOGrav 45-year and DSA-2000 25-year datasets instead, the upper limit on the local DM density in the form of compact objects can improve by more than two orders of magnitude compared to the published NANOGrav 15-yr dataset, reaching $f_{\mathrm{DM}}\lesssim 3$ for $M\gtrsim 10^{-5}$ $M_{\odot}$. We summarize the total observation time of each pulsar considered in the above analysis in Fig. 11.

### v2

@@ 4 Pulsar timing arrays
Millisecond pulsars (MSPs) have long been used as detectors of nanoHertz gravitational waves (GWs) [Taylor:2021yjx]. By measuring the timing of radio pulses from an array of MSPs accurately, one can measure ripples of spacetime manifesting as deviations in pulsar timings. In June 2023, The North American Nanohertz Observatory for Gravitational Waves (NANOGrav) announced positive evidence for a $\sim 3\sigma$ evidence of the existence of a stochastic gravitational wave background, possibly due to a population of merging supermassive BHs [NANOGrav:2023gor]. The DSA-2000 is projected to detect approximately 3,000 millisecond pulsars [2024AAS...24326104S], and to increase the total number of high-quality pulsars being timed to 200, with a monthly cadence [2019BAAS...51g.255H], significantly improving NANOGrav’s sensitivity to gravitational waves.
In addition to gravitational waves, PTAs have also been proposed as probes of DM substructure and primordial black holes (PBHs) [Siegel:2007fz, Seto:2007kj, Baghram:2011is, Kashiyama_2012, Clark:2015sha, Schutz:2016khr, Dror:2019twh, Kashiyama:2018gsh, Ramani:2020hdo, Gresham:2022biw, NANOGrav:2023hvm]. DM substructure are self-gravitating DM objects formed on small ($\lesssim$ kpc) scales. While the large-scale matter power spectrum is accurately measured from the cosmic microwave background (CMB), DM substructure on small scales is much more difficult to detect. More importantly, different DM theories can predict different small-scale structures [Hogan:1988mp, Kolb:1993zz, Erickcek:2011us, Fan:2014zua, Graham:2015rva], making the detection of these objects an important step in understanding the nature of DM. In this section, we review the idea of detecting DM substructure using PTAs, and explain how to calculate the projected sensitivity for a future PTA survey, including timing data acquired by DSA-2000. We find that the detection of DM substructure may be possible with sufficient observation time.
@@ 4.1 Gravitational effects
$\delta t(t)=\int_{0}^{t}dt^{\prime}\,\frac{\delta\Omega(t^{\prime})}{\Omega}\,$, (4.2)
where the timing residuals, $h(t)$, are defined to be the measured pulsar time shifts subtracting the best-fitted pulsar timing evolution consistent with the timing model of a naturally evolving pulsar [Taylor:2021yjx],
@@
$\frac{\delta\Omega(t)}{\Omega}=\int_{0}^{t}dt^{\prime}\,\mathbf{\hat{d}}\cdot\nabla\Phi_{\mathrm{grav}}(t^{\prime})\,$, (4.4)
where $\mathbf{d}$ is the vector along the line-of-sight between the pulsar and Earth. Combining Eqs. (4.1)-(4.4), one obtains the timing residual due to the Doppler effect from a transiting DM substructure [Dror:2019twh, Lee:2020wfn],
@@
Here we have absorbed all terms up to $\mathcal{O}(t^{2})$ into $\delta t_{\mathrm{fitted}}(t)$, since these terms are typically included in a fit within the timing model. Practically, when searching for physics signals within the PTA dataset, the fitting part is handled by software packages such as ENTERPRISE [2019ascl.soft12015E], ENTERPRISE_EXTENSIONS [enterprise], and a recently developed wrapper, PTArcade [Mitridate:2023oar]. See Ref. [Lee:2021zqw] for the Bayesian inference framework used to search for DM substructure with PTA data. Hence, the $\delta t_{\mathrm{fitted}}(t)$ term in Eq. (4.5) can generally be neglected when the signal shape is interfaced into these programs for searching the actual dataset. However, for the purpose of making sensitivity projections for future experiments, the $\delta t_{\mathrm{fitted}}(t)$ term can lead to a reduction in sensitivity if a portion of the signal is degenerate with the timing model and gets absorbed into the fit [Ramani:2020hdo].
Given a total observation time of $T$ on a particular pulsar, in the limit where $b\gg vT$ (which is applicable for DM masses sufficiently large), Eq. (4.5) can be expanded in a power series in tt. [^13] One can also consider signals in the opposite limit (i.e., the dynamic limit) and effects such as the Shapiro delay. However, their upper limits are generally weaker than that of the Doppler static effect when analyzing real data [NANOGrav:2023hvm], so we do not include them in this work. The terms up to quadratic order in $t$ are degenerate with the intrinsic pulsar timing model and will be absorbed into the fit. Therefore, we parameterize the timing residual starting from the cubic order in $t$ [Lee:2020wfn, Lee:2021zqw],
$\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}h(t)=\frac{A}{\mathrm{yr}^{2}}t^{3}-\delta t_{\mathrm{fitted}}(t)$}\,, (4.6)
@@
Here we have taken $t_{0}=0$ for simplicity. In addition, part of the signal in the cubic term of Eq. (4.7) will also be absorbed by the fit due to the finite observation time, $T$. This can be estimated by considering the general form of the quadratic timing model, $\delta t_{\mathrm{fit}}(t)=a_{0}+a_{1}t+\tfrac{1}{2}a_{2}t^{2}$, and solving for the coefficients $a_{0}$, $a_{1}$, and $a_{2}$ that minimize the difference, $\int_{0}^{T}dt\,\left[\frac{A}{\mathrm{yr}^{2}}t^{3}-\delta t_{\mathrm{fitted}}(t)\right]^{2}$. The resulting timing residual from Eq. (4.6) is given by [Ramani:2020hdo]
$\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}h(t)=\frac{A}{\mathrm{yr}^{2}}\left(t^{3}-\frac{3T}{2}t^{2}+\frac{3T^{2}}{5}t-\frac{T^{3}}{20}\right)$}\,. (4.8)
If the observed data contain a signal from DM, the measured timing residual, $d(t)$, can be written as the sum of contributions from noise and background, $n(t)$, and the DM signal
d(t)=n(t)+h(t).d(t)=n(t)+h(t)\,. (4.9)
The term $n(t)$ can generically be written as a sum of three distinct components
  • pulsar intrinsic white noise (frequency-independent),
  • pulsar intrinsic red noise (frequency-dependent) [2010ApJ...725.1607S],
  • a common spectrum red noise process (frequency-dependent) across all pulsars, possibly due to a stochastic gravitational wave background [1983ApJ...265L..39H].
For the majority of pulsars in NANOGrav’s 15-year data release, the common-spectrum process is the dominant contribution at low frequencies, while pulsar-intrinsic white noise dominates at high frequencies once the common-spectrum process power falls off sufficiently [NANOGrav:2023ctt]. Defining the one-sided noise power spectral density $S_{n}(f)$ as $\langle\tilde{n}(f)\tilde{n}^{*}(f^{\prime})\rangle=\tfrac{1}{2}S_{n}(f),\delta(f-f^{\prime})$ [Moore:2014lga] [^14] We follow the Fourier transform convention defined in Ref. [Moore:2014lga], namely $\tilde{x}(f)=\int_{0}^{T}dt\,x(t)\exp(-2\pi ift)$., we can parameterize $S_{n}(f)$ as
$S_{n}(f)=\left[N_{n}+\frac{A_{n}^{2}}{12\pi^{2}}\left(\frac{f}{\mathrm{yr}^{-1}}\right)^{-\gamma_{n}}\right]\mathrm{yr}^{3}\,$, (4.10)
where $N_{n}$, $A_{n}$, and $\gamma_{n}$ are dimensionless constants. Here we neglect spatial correlations of the common-spectrum process, since the DM pulsar-term signal under consideration is uncorrelated across pulsars. The theoretical prediction for the noise spectral index due to a stochastic gravitational wave background produced by merging super massive black hole binaries is $\gamma_{n}=13/3\approx 4.3$ [Phinney:2001di]. However, we note that the median and the 559595% quantiles of the posterior distribution reported in the NANOGrav 15-year analysis are $\gamma_{n}=3.2\pm 0.6$, with the best-fit value of $\gamma_{n}\approx 3.2$, assuming the Hellings–Downs spatial correlation [NANOGrav:2023gor].
The squared SNR can be estimated using the usual expression [Moore:2014lga]
$\mathrm{SNR}^{2}=4\int_{f_{\text{min}}}^{f_{\text{max}}}df\,\frac{|\tilde{h}(f)|^{2}}{S_{n}(f)}\,.$ (4.11)
For a finite observation time $T$ and cadence $\Delta t$ for a given pulsar, the time integral in the Fourier transform of the signal is cut off at $t=T$, while the frequency integral runs from $f_{\min}=1/T$ to the Nyquist frequency $f_{\max}=1/(2\Delta t)$. The (squared) Fourier transform of the DM signal can be readily computed from Eq. (4.8)
$|\tilde{h}(f)|^{2}$ $=\frac{A^{2}}{400\pi^{8}\mathrm{yr}^{4}}\frac{\left[\pi fT(\pi^{2}f^{2}T^{2}-15)\cos(\pi fT)-3(2\pi^{2}f^{2}T^{2}-5)\sin(\pi fT)\right]^{2}}{f^{8}}$
$\sim\frac{A^{2}}{400\pi^{2}\mathrm{yr}^{4}}\frac{T^{6}\cos^{2}(\pi fT)}{f^{2}}\,$, (4.12)
where we have taken $fT\gtrsim 1$ in the second line. Computing the SNR in Eq. (4.11) using the DM signal expression in Eq. (4.1) and the noise power spectral density in Eq. (4.10), we find [^15] We thank Abhiram Cherukupalli for pointing out the importance of including the red noise when deriving the scaling of the SNR with powers of $T$.
@@
$\mathrm{SNR}$ $\approx\frac{GMT^{3}v|\mathbf{\hat{v}}\cdot\mathbf{\hat{d}}|}{120\sqrt{2}\pi b^{3}\,\mathrm{yr}}\sqrt{\frac{1}{N_{n}}\left(\frac{\mathrm{yr}^{-1}}{f_{*}}\right)+\frac{12\pi^{2}}{(\gamma_{n}-1)A_{n}^{2}}\left(\frac{f_{*}}{\mathrm{yr}^{-1}}\right)^{\gamma_{n}-1}}$
$\equiv\frac{GMT^{3}v|\mathbf{\hat{v}}\cdot\mathbf{\hat{d}}|}{6b^{3}}\frac{1}{K_{n}}\,$, (4.13)
where we assume $\gamma_{n}>1$, define $f_{*}$ as the frequency where the white-noise and red-noise contributions are equal, and define the quantity $K_{n}$ (with dimensions of [time]) as the effective noise level that enters the SNR. Here we note that the SNR computed in Eq. (4.1) is dominated by the contribution at the noise turnover frequency, $f_{*}$. This can be easily understood by noting that $|\tilde{h}(f)|^{2}\sim f^{-2}$ for $f\gtrsim 1/(2\pi T)$ as in Eq. (4.1), and that the noise scales as $S_{n}(f)\propto f^{-\gamma_{n}}$ for $f\lesssim f_{*}$, and $S_{n}(f)\sim\mathrm{constant}$ for $f\gtrsim f_{*}$. Therefore, the integrand in Eq. (4.11) scales as $f^{\gamma_{n}-2}$ and $f^{-2}$ at low and high frequencies, respectively, and thus as long as $\gamma_{n}>1$, the SNR integral will be dominated by frequency near $f=f_{*}$.
@@
The above analysis considers the SNR from timing measurements of a single pulsar. We now generalize this to the case of PTAs, where timing measurements from multiple pulsars are available. Since the DM phase space is a statistical distribution near each pulsar, the expected SNR from each pulsar also follows a statistical distribution. The variance in DM velocity can generally be neglected in an order-of-magnitude estimate due to its linear scaling in the SNR. Statistically, the signal is dominated by the DM with the smallest impact parameter. For a given pulsar, this can be derived by considering putting random points in a spherical volume, in which the cumulative distribution function (CDF) of the impact parameter of each DM is given by $F_{B}(b)=(b/R)^{3}$, where $R$ is the radius of a large (fictitious) spherical volume, and thus the CDF of the minimal impact parameter is [Dror:2019twh]
@@
$={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\exp\left(-\frac{4\pi}{3}\frac{G\rho_{\mathrm{DM}}f_{\mathrm{DM}}}{6}\frac{1}{\mathrm{SNR}}\frac{T^{3}v|\mathbf{\hat{v}}\cdot\mathbf{\hat{d}}|}{K_{n}}\right)}\,.$ (4.15)
@@
to be the fraction of dark matter in the form of compact objects, which are the components that can contribute to PTA signals. The theoretically compelling parameter space is $f_{\mathrm{DM}}\leq 1$, although this assumes the robustness of the canonical local dark matter density value, $\rho_{\mathrm{DM}}=0.4~\mathrm{GeV/cm}^{3}$. Current experimental constraints, however, allow values as large as $f_{\mathrm{DM}}\lesssim 2\times 10^{3}$, based on limits from solar system ephemerides [Pitjev_2013], since there is currently no direct local observational measurement of $\rho_{\mathrm{DM}}$.
If we are now given timing data from $N_{P}$ pulsars, each labeled with subscript $i$, the optimal sensitivity to compact DM is achieved by considering the statistics of the pulsar with the highest SNR [Lee:2021zqw], denoted as $\mathrm{SNR}_{\max}$, whose CDF can be readily derived from Eq. (4.1), and the fact that the SNRi from each pulsar is independent:
@@
$={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\exp\left(-\frac{4\pi}{3}\frac{G\rho_{\mathrm{DM}}f_{\mathrm{DM}}}{6}\frac{1}{\mathrm{SNR}_{\max}}\sum_{i=1}^{N_{P}}\frac{T_{i}^{3}v_{i}|\mathbf{\hat{v}}_{i}\cdot\mathbf{\hat{d}}_{i}|}{K_{n,i}}\right)}\,$, (4.17)
@@
where $K_{n,i}$ is the effective noise level as defined in Eq. (4.1) for each pulsar. Using Eq. (4.1), for any $0\leq p\leq 1$, we can derive the $100p^{\mathrm{th}}$-percentile $\mathrm{SNR}_{\max}$, denoted as $\mathrm{SNR}_{\max,p}$, by setting $F_{\mathrm{SNR}_{\max}}(\mathrm{SNR}_{\max,p})=p$,
$\mathrm{SNR}_{\max,p}$ $={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\frac{4\pi}{3}\frac{G\rho_{\mathrm{DM}}f_{\mathrm{DM}}}{6}\frac{1}{\log p^{-1}}\sum_{i=1}^{N_{P}}\frac{T_{i}^{3}v_{i}|\mathbf{\hat{v}}_{i}\cdot\mathbf{\hat{d}}_{i}|}{K_{n,i}}}$
$\propto f_{\mathrm{DM}}\sum_{i=1}^{N_{P}}T_{i}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}3}}\,$, (4.18)
@@
where $N_{P}$ is the number of pulsars used in the analysis. We have assumed that the effective noise level for timing data from each pulsar is comparable, and that the noise level for new pulsars to be discovered and timed by DSA-2000 is the same as that of existing pulsars timed by NANOGrav. Additionally, in the second line, we have neglected the variance in DM velocity and the variance $\mathcal{O}(1)$ angular factors Note in particular that the pulsar observing time $T_{i}$ has a particularly strong impact on the SNR, since a longer observation time leads to a smaller expectation value of $b_{\min}$, and thus a stronger signal.
@@
Figure 10: Upper limits at the 95% credible level on the local PBH abundance (left panel) and fifth force coupling strength (right panel). The blue line is the existing constraint published by the NANOGrav collaboration [NANOGrav:2023hvm], while the orange and green lines are the projected sensitivity by 2033 and 2053, corresponding to 5 and 25 years after DSA-2000 begins taking timing data, obtained by rescaling the blue line using the projected parameters for a combined NANOGrav 25-yr / DSA 5-yr and NANOGrav 45-yr / DSA 25-yr observation dataset, and Eq. (4.19) and Eq. (4.22) (see main text). Following Ref. [NANOGrav:2023hvm], existing constraints on $\tilde{\alpha}$ are also shown in the right panel, including bounds from the weak equivalence principle (WEP) [Wagner:2012ui, Shao:2018klg, Sun:2019ico], NS kinetic heating (assuming an additional short-range DM–baryon interaction) due to DM capture [Gresham:2022biw], and bounds derived from a combination of fifth-force constraints on baryon–baryon interactions [Berge:2017ovy, PhysRevD.99.055043] and Bullet Cluster (BC) limits on DM self-interactions [Spergel:1999mh, Kahlhoefer:2013dca] (see also Refs. [Coskuner:2018are, Gresham:2022biw]). We note that while the BC bound is very competitive, it is an indirect constraint on $\tilde{\alpha}$ and does not apply if only a subcomponent of DM interacts through the fifth force, where, in contrast, the other constraints degrade only linearly with the DM subcomponent fraction.
@@
Figure 11: Observation times for each pulsar considered in the analysis are shown for three benchmark years: 2023 (blue), 2033 (orange), and 2053 (green). The 68 existing pulsars being timed by NANOGrav in their 15-year data release (2023) are shown at the top [NANOGrav:2023hde], although J0614-3329 (marked by an asterisk) is excluded from their DM substructure analysis since its observation time has been less than three years [NANOGrav:2023hvm]. The new pulsars that DSA-2000 is expected to discover are shown at the bottom, labeled with placeholder pulsar names starting with “New Pulsar”.
@@
$\frac{f_{\mathrm{DM}}}{f_{\mathrm{DM}}^{(\mathrm{NG-15})}}=\frac{\sum_{j}^{N_{P}^{\mathrm{NG-15}}}(T^{(\mathrm{NG-15})}_{j})^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}3}}}{\sum_{i}^{N_{P}}T_{i}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}3}}}\,$, (4.19)
where $N_{P}^{(\mathrm{NG-15})}=67$ is the number of pulsars used in the NANOGrav 15-yr new physics analysis (one pulsar, J0614-3329, was excluded from the analysis since it was observed for less than three years) [NANOGrav:2023hvm], released in 2023.
The above analysis shows that the sensitivity to DM substructure can be improved by (i) obtaining timing data from more pulsars, and (ii) a longer cumulative observation time for each pulsar. Both of these conditions can be achieved by utilizing DSA-2000’s power in discovering new pulsars. DSA-2000 is expected to bring the total number of pulsars being timed within the NANOGrav Collaboration to 200. In the left panel of Fig. 10, we show the projected sensitivity from the outlined procedure, assuming that the new pulsars start being timed by the collaboration in 2028. We show the projected sensitivity by 2033 and 2053, corresponding to 5 and 25 years after DSA-2000 begins taking timing data. We find that with the combined data from the NANOGrav 25-yr and DSA-2000 5-year datasets, which is the currently planned timeframe for DSA-2000, the upper limit that one can place on the local DM density in the form of compact objects can improve by roughly an order of magnitude, to $f_{\mathrm{DM}}\lesssim{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}50}$ for $\gtrsim 10^{-5} MM_{\odot}$. If we consider the combined data from the NANOGrav 45-year and DSA-2000 25-year datasets instead, the upper limit on the local DM density in the form of compact objects can improve by more than two orders of magnitude compared to the published NANOGrav 15-yr dataset, reaching $f_{\mathrm{DM}}\lesssim{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}4}$ for $\gtrsim 10^{-5} MM_{\odot}$. We summarize the total observation time of each pulsar considered in the above analysis in Fig. 11.